OpenStudy (anonymous):

If 1/a +1/b = 10, what is the value of a+b ?

OpenStudy (anonymous):

a + b = 10ab

OpenStudy (anonymous):

Start with the original equation and multiply everything by ab This leaves \[ab(\frac{1}{a}) + ab(\frac{1}{b}) = ab(10)\] \[\frac{ab}{a} + \frac{ab}{b} = 10ab\] \[b + a = 10ab\]

OpenStudy (anonymous):

You can test it with various values of a and b to see if it works. I checked a few and it works.

OpenStudy (anonymous):

Oh ok thank you